OpenStudy (anonymous):

2(12+-8x)=1x+-11x need help solving

OpenStudy (texaschic101):

2(12 + (-8x) = x + (-11x) 2(12 - 8x) = x - 11x 24 - 16x = -10x 24 = -10x + 16x 24 = 6x 4 = x check... 2(12 - 8(4)) = 4 - 11(4) 2(12 - 32) = 4 - 44 2(- 20) = -40 -40 = -40 (correct) so x = 4 any questions ?